OpenStudy (the_puppet1):

Solve the equation -21 + 25n = 14 for n. Help!!!

OpenStudy (erak):

Move the "-21" over to the right side, then solve for n. Try it

OpenStudy (sagewilson):

In order to find N you'll need to add 21 to both sides, this will remove the -21. You'll be left with 25N = 14 + 21 This will give you what?

RhondaSommer (rhondasommer):

what do you have

OpenStudy (dcolley):

What do you think the answer is?

RhondaSommer (rhondasommer):

add 21 to both sides of the equal sign \[−21+25n=14\]\[+21~~~~~~~~~~~~+21\]\[−−−−−−−−\]

OpenStudy (the_puppet1):

21 + 21 = 42

RhondaSommer (rhondasommer):

.-. no like -21+21=? then 14+21=?

OpenStudy (the_puppet1):

Oh :P

RhondaSommer (rhondasommer):

Ill write it differently. \[-21+21+25n=14+21\]

OpenStudy (sagewilson):

So -21 + 21 is..?

RhondaSommer (rhondasommer):

u still there?

OpenStudy (the_puppet1):

0

RhondaSommer (rhondasommer):

yep. 14+21=?

OpenStudy (the_puppet1):

35

RhondaSommer (rhondasommer):

now you divide both sides by 25\[\frac{25n}{25}=\frac{35}{25}\]

RhondaSommer (rhondasommer):

\[\frac{25n}{25}=?\]

RhondaSommer (rhondasommer):

\[\frac{35}{25}=?\]

OpenStudy (the_puppet1):

It's A cause 35 Divided by 25 = 1/4

RhondaSommer (rhondasommer):

yup :D

RhondaSommer (rhondasommer):

no. no it isnt.

RhondaSommer (rhondasommer):

its 1.4 not 1/4

RhondaSommer (rhondasommer):

when you divide all the fractions that are available as answers and see which one equals 1.4

OpenStudy (the_puppet1):

@RhondaSommer

OpenStudy (the_puppet1):

Idk if the 2nd is correct: Solve the equation 14 + 14p = 21 for p. A) 1/2 B) 1/3 C) 1/6 D) 1/8 I think it's D but I'm not sure if I'm correct

OpenStudy (carolina4567):

Its not D , give it another try

OpenStudy (griefer101):